OpenStudy (anonymous):

find the indicated function composition and evaluate. f(g(x)) and g(f(x)) f(x)=2x^2+2 g(x)=6x-5

OpenStudy (gorv):

f(g(x))=f(6x-5)=2*(6x-5)^2+2

OpenStudy (gorv):

g(f(x))=g(2x^2+2)=6*(2x^2+2)-5